Consider two displacements, one of magnitude 15 m and another of magnitude 20 m. What angle between the directions of this two displacements give a resultant displacement of magnitude (a) 35 m, (b) 5 m, and (c) 25 m.
Answer #1

using the equation

R^2 = F1^2 + F2^2 + 2F1F2 cos θ

a) 35^2 = 15^2 + 20^2 + 2*15*20*cos θ

θ = 0o

b) 5^2 = 15^2 + 20^2 + 2*15*20*cos θ

θ = 180o

c) 25^2 = 15^2 + 20^2 + 2*15*20*cos θ

θ = 90o